# The functional Loomis-Whitney type inequality in the Heisenberg groups and Projection theorems over finite fields

### v1

@@ 2.1 Basics on subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
By a group structure on a vector space $V$ over $\mathbb{F}_{q}$ we mean the underlying additive group structure on $V$, forgetting the scalar multiplication. Trivially, a subspace of the vector space $V$ is a subgroup of $V$. If the field $\mathbb{F}_{q}$ is a prime field, i.e., $q=p$ for some prime $p$, then the converse is also true. Indeed, if $W$ is a subgroup of $V,$ then we can give a scalar multiplication on $W$ that makes $W$ into a subspace of $V.$ To see this, we may assume $\mathbb{F}_{p}=\{0,1,\ldots,p-1\}.$ For $t\in\mathbb{F}_{p}$ and $w\in W$, define $tw:=w+\cdots+w$ ($t$-times). We easily see that this coincides with the scalar multiplication on $W$ induced from $V$, and so $W$ is a subspace of $V.$ If $\mathbb{F}_{q}$ is not a prime field, then there is a subgroup $W$ of $V$ that is not a subspace. For example, if $q=p^{r}$ for some $r\geq 2,$ the subfield $\mathbb{F}_{p}$ is a subgroup of the (one dimensional) vector space $\mathbb{F}_{q},$ but not a subspace of $\mathbb{F}_{q}$ (over $\mathbb{F}_{q}$). Over a general $\mathbb{F}_{q}$ with $q=p^{r},$ in order to obtain all subgroups of a vector space $V$, it is enough to view $V$ as a vector space over $\mathbb{F}_{p}$ and find all subspaces of $V$ over $\mathbb{F}_{p}.$ Similarly, linear maps between two vector spaces over a prime field are precisely group homomorphisms between the corresponding underlying groups.
@@
Define an action of the group $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$ by $s\cdot(x,t)=(sx,s^{2}t)$ for $s\in\mathbb{F}_{q}^{*}$ and $(x,y)\in\mathbb{H}^{n}.$ This action shall be called the dilation action, following [2].
@@
A subgroup $G$ of $\mathbb{H}^{n}$ is called homogeneous if $G$ is invariant under the dilation action of $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}$.
Recall that the orbit of an element $(x,t)\in\mathbb{H}^{n}$ under the dilation action is given by
@@
If $G$ is any homogeneous subgroup of $\mathbb{H}^{n}$ containing $(x,t)$, then $G$ necessarily includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$.
@@
We can easily check that the isotropy groups of $(x,t)$ for the cases (i), (ii), (iii)
@@
are $\{0\},$ $\mathbb{Z}_{2}$ and $\mathbb{F}_{q}^{*}$ , respectively. Then the lemma is immediate from the formula
$|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)||\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)|=|\mathbb{F}_{q}^{*}|=q-1.$
@@
$\Gamma_{S}(\rho):=\{(x,t)\subset\mathbb{H}^{n}\,|\,x\in S,t=\rho(x)\}.$
@@
Next, let us treat homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G$ is a homogeneous subgroup of $\mathbb{H}^{n},$ then $\pi_{h}(G)$ is a subspace of $\mathbb{F}_{q}^{2n}$ (over $\mathbb{F}_{q}$). So fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$ and find all subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S.$ Assume that $q=p^{r}$ for some prime $p$ and positive integer $r.$ The projection $\pi_{h}:\mathbb{H}^{n}\rightarrow\mathbb{F}_{q}^{2n}$ induces a group homomorphism $\pi_{h}|_{G}:$ $G\rightarrow S.$ Then elements of the kernel $K$ of $\pi_{h}|_{G}$ are of the form $(0,t)$ for $t\in\mathbb{F}_{q}$. Also, note that the restriction $\pi_{v}|_{K}:K\rightarrow\mathbb{F}_{q}$ embeds $K$ into $\mathbb{F}_{q}$ as a subgroup. Thus $|K|$ divides $q=p^{r}$, and so $|K|=p^{\ell}$ for some $\ell\leq r.$ Firstly, assume that $K$ is nontrivial, i.e., contains an element $(0,t)$ with $t\neq 0.$ Since $K$ is homogeneous, $K$ includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(0,t)$. So, by Lemma 2.3 (ii), we obtain $|K|\geq\frac{q-1}{2}=\frac{p^{r}-1}{2}$. This forces $|K|=p^{r}=q$ since $p$ is odd. (Note that if $\ell\leq r-1$, then it would be that $|K|<\frac{q-1}{2}.$) Since $|G|=|S||K|=|S|q,$ in this case we have $G=S\times\mathbb{F}_{q}.$ Secondly, assume that $K=\{(0,0)\}$ is trivial. Then $\pi_{h}|_{G}:G\rightarrow S$ is an isomorphism, and so, as in Proposition 2.4, $S$ is isotropic and $G=\Gamma_{S}(\rho)$ for some group homomorphism(equivalently, a linear map over $\mathbb{F}_{p}$) $\rho:S\rightarrow\mathbb{F}_{q}.$ We claim that $\rho$ is trivial. To show this, let $(x,\rho(x))\in\Gamma_{S}(\rho)=G$ with $0\neq x\in\mathbb{F}_{q}^{2n}$. Then by the homogeneity condition, for any $s\in\mathbb{F}_{q}^{*}$, we have
@@
But, in particular, if $s\in\mathbb{F}_{p}^{*}$, then $\rho(sx)=s\rho(x)$ since $\rho$ is linear over $\mathbb{F}_{p}.$ By this fact and (16), we obtain the equality $s^{2}\rho(x)=s\rho(x)$ for any $s\in\mathbb{F}_{p}^{*}$. This implies that $\rho(x)=0$ for all nonzero $x\in\mathbb{F}_{q}^{2n}$, i.e., $\rho$ is trivial. Therefore, $G=\Gamma_{S}(\rho)=S\times\{0\}.$ In conclusion, we have proved:
@@ 2.2 Orthogonal complement of a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$
For a subspace $V$ of the vector space $\mathbb{F}_{q}^{m},$ let $V^{\perp}$ be the subspace of $V$ defined as $V^{\perp}=\{w\in\mathbb{F}_{q}^{m}:\langle v,w\rangle=0\,\,\mathrm{for\,\,\,all}\,v\in V\}$, where $\langle\,$$\rangle$ is the standard inner product on $\mathbb{F}_{q}^{m}.$ We may call $V^{\perp}$ the orthogonal complement of $V$ w.r.t. $\langle\,$$\rangle$. By Proposition 2.5, homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ are some subspaces of $\mathbb{F}_{q}^{2n+1},$ and so we can take $G^{\perp}$ for homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\{0\}$ for an isotropic $S$ is a homogeneous subgroup, then we have $G^{\perp}=S^{\perp}\times\mathbb{F}_{q}$, which is necessarily a (vertical) homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\mathbb{F}_{q}$, where $S$ is an arbitrary subspace of $\mathbb{F}_{q}^{2n}$, then $G^{\perp}=S^{\perp}\times\{0\}$. In this case $G^{\perp}$ is not necessarily a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, if $\mathrm{dim}\,S\leq n-1$, then $\mathrm{dim}\,S^{\perp}\geq n+1$ and so $S$ is not isotropic by the dimension reason. Thus $G^{\perp}=S^{\perp}\times\{0\}$ is not a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. (In fact, it is not even a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$.)
@@ 2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
Now we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for the two special cases in the above. For $k=0,\ldots,2n$, let $\mathrm{Gr}(k,2n)$ be the variety of $k$ dimensional subspaces in $\mathbb{F}_{q}^{2n}.$ For $k=0,\ldots n,$ let $\mathrm{IG}(k,2n)$ be the variety of $k$ dimensional isotropic subspaces in $\mathbb{F}_{q}^{2n}$.
First we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, using Proposition 2.4. For a subspace $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let n(S)n(S)(resp. $n(S)^{\prime}$) be the number of subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then, by Proposition 2.4, the number of all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ equals
@@
We easily see that $n(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $n(S)=kp$ for each $S\in\mathrm{IG}(k,2n)$ because there are $kp$ group homomorphisms (linear maps) from $S$ to $\mathbb{F}_{p}$, identified with $1\times k$ matrices with entries in $\mathbb{F}_{p}$. This number equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}kp|\mathrm{IG}(k,2n)|.$ (17)
Next, we count homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$, using Proposition 2.5. For each subspaces $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let m(S)m(S)(resp. $m(S)^{\prime}$) be the number of homogeneous subgroups $G$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then by Proposition 2.5, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
@@
Then the following formulas can be found in the literature, for example, in [22, 5] .
@@
For $1\leq r,s\leq\infty,$ denote by $A(s\to r)$ the smallest number, if it exist, such that the inequality
@@
For $1\leq u_{1},u_{2}\leq\infty,$ define $L(u_{1},u_{2})$ as the smallest positive constant, if it exist, such that the following estimate holds for all nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$
@@
and if it does not exist, we set $L(u_{1},u_{2}):=\infty.$
Recall that for $1\leq u\leq\infty$, the Holder’s conjugate of $u$ is defined by $\frac{1}{u}+\frac{1}{u^{\prime}}=1$.
@@
Since these inequalities holds for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, the lemma is immediate. ∎
@@
But these estimates follows directly from Lemma 3.2 and the following result, due to Koh [16, Theorem 1.1]:
@@
To prove the lemma, it is sufficient to show the LHS of the inequality (7) with $n=1$ is equal to LHS, $\mathcal{L}(f_{1},f_{2})$, of the inequality (20). In fact, it follows from (1) that
@@
where the last equality we have used the fact that $f_{n}$-term is independent of $x_{n}$. By applying Hölder inequality with the conjugate pair of exponents $(2n,\frac{2n}{2n-1})$, we obtain
@@
Therefore it remains to bound
@@
$f_{j}(\pi_{j}(x^{\prime},y,\tau))=\begin{cases}f_{j}(\check{x}_{j},\tau+\frac{1}{2}x_{j}x_{j+n}),~~~\text{ if }~~~~~~1\leq j\leq n-1,\\ f_{j}(\check{x}_{n+j},\tau-\frac{1}{2}x_{j-n}x_{j}),~~~\text{ if }n+1\leq j\leq 2n-1.\end{cases}$ (24)
@@
$J_{\Pi}\leq$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}$
@@
$\times\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}^{2n-1}\bigg)^{\frac{1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
@@
$I=$ $\frac{1}{q^{2n+1}}=\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
@@
$\leq$ $\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\cdot\|f_{2n}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}^{2n}_{q},dx)}.$ (29)

### v2

@@ 2.1 Basics on subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
When viewing a vector space $V$ over $\mathbb{F}_{q}$ as a group, we consider only its additive group structure, forgetting scalar multiplication. Trivially, a subspace of the vector space $V$ is a subgroup of $V$. If the field $\mathbb{F}_{q}$ is a prime field, i.e., $q=p$ for some prime $p$, then the converse is also true. Indeed, if $W$ is a subgroup of $V,$ then we can give a scalar multiplication on $W$ that makes $W$ into a subspace of $V.$ To see this, we write $\mathbb{F}_{p}=\{0,1,\ldots,p-1\}.$ For $t\in\mathbb{F}_{p}$ and $w\in W$, define $tw:=w+\cdots+w$ ($t$-times). We easily see that this coincides with the scalar multiplication on $W$ induced from $V$, and so $W$ is a subspace of $V.$ If $\mathbb{F}_{q}$ is not a prime field, then there is a subgroup $W$ of $V$ that is not a subspace. For example, if $q=p^{r}$ for some $r\geq 2,$ the subfield $\mathbb{F}_{p}$ is a subgroup of the (one dimensional) vector space $\mathbb{F}_{q},$ but not a subspace of $\mathbb{F}_{q}$ (over $\mathbb{F}_{q}$). Over a general $\mathbb{F}_{q}$ with $q=p^{r},$ in order to obtain all subgroups of a vector space $V$, it is enough to view $V$ as a vector space over $\mathbb{F}_{p}$ and find all subspaces of $V$ over $\mathbb{F}_{p}.$ Similarly, linear maps between two vector spaces over a prime field are precisely group homomorphisms between the corresponding underlying groups. For a subspace $W$ of the vector space $\mathbb{F}_{q}^{m},$ let $W^{\perp}$ be the subspace of $\mathbb{F}_{q}^{m}$ defined as $W^{\perp}=\{v\in\mathbb{F}_{q}^{m}:\langle v,w\rangle=0\,\,\mathrm{for\,\,\,all}\,w\in W\}$, where $\langle\,$$\rangle$ is the standard inner product on $\mathbb{F}_{q}^{m}.$ We may call $W^{\perp}$ the orthogonal complement of $W$ in $\mathbb{F}_{q}^{m}$ with respect to $\langle\,$$\rangle$.
@@
Define an action of the group $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$ by $s\cdot(x,t)=(sx,s^{2}t)$ for $s\in\mathbb{F}_{q}^{*}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}).$ This action shall be called the dilation action, following [2].
@@
A subgroup $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ is called homogeneous if $G$ is invariant under the dilation action of $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$.
Recall that the orbit of an element $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$ under the dilation action is given by
@@
If $G$ is any homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$ containing $(x,t)$, then $G$ necessarily includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$.
@@
It can be easily verified that the isotropy groups of $(x,t)$ for the cases (i), (ii), (iii)
@@
are $\{1\},$ $\mathbb{Z}_{2}=\{-1,1\}$ and $\mathbb{F}_{q}^{*}$ , respectively. Then, the lemma is immediate from the formula
$|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|\cdot|\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)|=|\mathbb{F}_{q}^{*}|=q-1.$
@@
$\Gamma_{S}(\rho):=\{(x,t)\subset\mathbb{H}^{n}(\mathbb{F}_{p})\,|\,x\in S,t=\rho(x)\}.$
@@
Next, let us treat homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G$ is a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then $\pi_{h}(G)$ is a subspace of $\mathbb{F}_{q}^{2n}$ (over $\mathbb{F}_{q}$). So, we fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$, then find all subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S.$ Assume that $q=p^{r}$ for some prime $p$ and a positive integer $r$. The projection $\pi_{h}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}^{2n}$ induces a group homomorphism $\pi_{h}|_{G}:$ $G\rightarrow S.$ Then, elements of the kernel $K$ of $\pi_{h}|_{G}$ are of the form $(0,t)$ for $t\in\mathbb{F}_{q}$. Also, note that the restriction $\pi_{v}|_{K}:K\rightarrow\mathbb{F}_{q}$ embeds $K$ into $\mathbb{F}_{q}$ as a subgroup. Thus, $|K|$ divides $q=p^{r}$, and so $|K|=p^{\ell}$ for some $\ell\leq r.$ Firstly, assume that $K$ is nontrivial, i.e., contains an element $(0,t)$ with $t\neq 0.$ Since $K$ is homogeneous, $K$ includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(0,t)$. So, by Lemma 2.3 (ii), we obtain $|K|\geq\frac{q-1}{2}=\frac{p^{r}-1}{2}$. This forces $|K|=p^{r}=q$ since $p$ is odd. (Note that if $\ell\leq r-1$, then it would be that $|K|<\frac{q-1}{2}.$) Since $|G|=|S||K|=|S|q,$ in this case we have $G=S\times\mathbb{F}_{q}.$ Secondly, assume that $K=\{(0,0)\}$ is trivial. Then $\pi_{h}|_{G}:G\rightarrow S$ is an isomorphism, and so, as in Proposition 2.4, $S$ is isotropic and $G=\Gamma_{S}(\rho)$ for some group homomorphism(equivalently, a linear map over $\mathbb{F}_{p}$) $\rho:S\rightarrow\mathbb{F}_{q}.$ We claim that $\rho$ is trivial. To show this, let $(x,\rho(x))\in\Gamma_{S}(\rho)=G$ with $0\neq x\in\mathbb{F}_{q}^{2n}$. Then by the homogeneity condition, for any $s\in\mathbb{F}_{q}^{*}$, we have
@@
But, in particular, if $s\in\mathbb{F}_{p}^{*}$, then $\rho(sx)=s\rho(x)$ since $\rho$ is linear over $\mathbb{F}_{p}.$ By this fact and (12), we obtain the equality $s^{2}\rho(x)=s\rho(x)$ for any $s\in\mathbb{F}_{p}^{*}$. This implies that $\rho(x)=0$ for all nonzero $x\in\mathbb{F}_{q}^{2n}$, i.e., $\rho$ is trivial. Therefore, $G=\Gamma_{S}(\rho)=S\times\{0\}.$ In conclusion, we have proved
@@ 2.2 Orthogonal complement of a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$
By Proposition 2.5, homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ are some subspaces of $\mathbb{F}_{q}^{2n+1}$ via the identification $\mathbb{H}^{n}(\mathbb{F}_{q})=\mathbb{F}_{q}^{2n+1}$. By abuse of notation, through this identification, we take $G^{\perp}$ for homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\{0\}$ for an isotropic $S$ is a homogeneous subgroup, then we have $G^{\perp}=S^{\perp}\times\mathbb{F}_{q}$, which is necessarily a (vertical) homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\mathbb{F}_{q}$, where $S$ is an arbitrary subspace of $\mathbb{F}_{q}^{2n}$, then $G^{\perp}=S^{\perp}\times\{0\}$. In this case $G^{\perp}$ is not necessarily a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, if $\mathrm{dim}\,S\leq n-1$, then $\mathrm{dim}\,S^{\perp}\geq n+1$ and so $S$ is not isotropic by the dimension reason. Thus $G^{\perp}=S^{\perp}\times\{0\}$ is not a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. (In fact, it is not even a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$.)
@@ 2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
Now we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for the two special cases in the above. For $k=0,\ldots,2n$, let $\mathrm{Gr}(k,2n)$ be the variety of $k$ dimensional subspaces in $\mathbb{F}_{q}^{2n}.$ For $k=0,\ldots,n$, let $\mathrm{IG}(k,2n)$ be the variety of $k$ dimensional isotropic subspaces in $\mathbb{F}_{q}^{2n}$.
First we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, using Proposition 2.4. For a subspace $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let n(S)n(S)(resp. $n(S)^{\prime}$) be the number of subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{p})$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then, by Proposition 2.4, the number of all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ equals
@@
We easily see that $n(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $n(S)=p^{k}$ for each $S\in\mathrm{IG}(k,2n)$ because there are $p^{k}$ group homomorphisms (linear maps) from $S$ to $\mathbb{F}_{p}$, identified with $1\times k$ matrices with entries in $\mathbb{F}_{p}$.
This number equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}p^{k}|\mathrm{IG}(k,2n)|.$ (13)
Next, we count homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$, using Proposition 2.5. Fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$ of dimension $k$ (so $|S|=q^{k}$). For such $S$, let $m(S)$ (resp. $m(S)^{\prime}$) be the number of homogeneous subgroups $G$ with $\pi_{h}(G)=S$ and $|G|=q^{k}$ (resp. $|G|=q^{k+1}$). Then, by Proposition 2.5, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
@@
Then, the following formulas can be found in the literature, for example, in [5, 19].
@@
In conclusion, the coordinate subgroups $W_{1},\dots,W_{2n}$ discussed here form a homogeneous family that will be central to our main results. A detailed analysis of other homogeneous subgroups will be reserved for future work, to maintain focus and clarity in the current discussion.
@@
For $1\leq r,s\leq\infty,$ denote by $A(s\to r)$ the smallest number, if it exists, such that the inequality
@@
For $1\leq u_{1},u_{2}\leq\infty,$ define $L(u_{1},u_{2})$ as the smallest positive constant, if it exists, such that the following estimate holds for all nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$
@@
and if it does not exist, we set $L(u_{1},u_{2})=\infty.$
Recall that for $1\leq u\leq\infty$, the Hölder’s conjugate of $u$ is defined by $\frac{1}{u}+\frac{1}{u^{\prime}}=1$.
@@
Since these inequalities hold for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, the lemma is immediate. ∎
@@
But these estimates follow directly from Lemma 3.2 and the following result, due to Koh [13, Theorem 1.1]:
@@
To prove the lemma, it is sufficient to show the LHS of the inequality (7) with $n=1$ is equal to LHS, $\mathcal{L}(f_{1},f_{2})$, of the inequality (16). In fact, it follows from (7.1) that
@@
where in the last equality we have used the fact that $f_{n}$-term is independent of $x_{n}$. By applying Hölder inequality with the conjugate pair of exponents $(2n,\frac{2n}{2n-1})$, we obtain
@@
Therefore, it remains to bound
@@
To pass from dimension $n$ to n1n-1, we isolate the last pair of coordinates $(x_{n},x_{2n})$ and treat it as a parameter. For each fixed choice of $(x_{n},x_{2n})\in\mathbb{F}_{q}^{2}$, the remaining variables (together with the $t$-variable after the changes made in Steps $1^{\circ}$$2^{\circ}$) carry the same Heisenberg structure as $\mathbb{H}^{\,n-1}(\mathbb{F}_{q})$. With this preparation, the bracketed expression in (4) can be viewed as an (n1)(n-1)–dimensional Loomis–Whitney type form in the remaining variables, applied to auxiliary factors obtained from the original $f_{j}$’s by freezing $(x_{n},x_{2n})$. This allows us to invoke the induction hypothesis and obtain (4). The specific power $\frac{2n+1}{2n-1}$ is chosen so that the norms produced by the (n1)(n-1)–dimensional estimate translate, after averaging over $(x_{n},x_{2n})$, into the desired $L^{2n+1}$ and $L^{\frac{2n+1}{2}}$ norms of the original functions.
@@
$f_{j}(\pi_{j}(x^{\prime},y,\tau))=\begin{cases}f_{j}(\check{x}_{j},\tau+\frac{1}{2}x_{j}x_{j+n}),~~~~~~~~~~~~~\text{ if }1\leq j\leq n-1,\\ f_{j}(\check{x}_{n+j},\tau-\frac{1}{2}x_{j-n}x_{j}),~~~\text{ if }n+1\leq j\leq 2n-1.\end{cases}$ (20)
@@
$J_{\Pi}\lesssim$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}$
@@
$\cdot\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}^{2n-1}\bigg)^{\frac{1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
@@
$I=$ $\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
@@
$\leq$ $\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\cdot\|f_{2n}\|_{L^{\ell}(\mathbb{F}^{2n}_{q},dx)}.$ (25)